A solution of magnesium sulfate is 0.0500 molal. Calculate the mass percentage of MgSO4 in this solution.


1
Expert's answer
2020-12-10T07:52:26-0500

Mass % = [mass of solute/mass of solution]100%

Assume 100g of a solution

Let mass %= X%

Mass of solute= [mass%/100] mass of solution

=[X%/100%]100g

=Xg MgSO4

Moles of MgSO4=X/120.366

                       =0.008308Xmoles

Mass of solution=mass of solute mass of solvent

X% MgSO4 by mass=100g solution

Xg MgSO4= 0.008308Xmoles MgSO4

Mass of solvent=Mass of solution- Mass of solute

100g-Xg=[100-X] g=[100-X]/1000Kg of solvent

0.0500=[0.008308X]1000/(100-X)

[100-x]0.05=8.308X

5-0.05X=8.308X

X=5/8.358

X=0.5982g MgSO4

Mass %= [0.5982/100g]100%

=0.5982%